RhondaSommer (rhondasommer):

IM BACK :) Given 🔺ABC with a = 24, angle A = 108⁰ and angle B = 33⁰, find b. Round the sine value to the nearest thousandth and the answer to the nearest hundredth. 39.00 17.83 15.88 13.75

RhondaSommer (rhondasommer):



RhondaSommer (rhondasommer):

@sweetburger :) Sean can you help

OpenStudy (sweetburger):

This question we can use the law of sins. \[\sin a /a = \sin b /b\] So we can write this as

OpenStudy (sweetburger):

s\[\sin(108)/ 24 =\sin(33)/b\]

OpenStudy (sweetburger):

we can rewrite this to look like b *sin(108) = 24*sin(33)

OpenStudy (sweetburger):

divide both sides by sin(108) and you will have b by itself

RhondaSommer (rhondasommer):

couldnt you just cross multiply?

OpenStudy (sweetburger):

Yes, that is what I did.

RhondaSommer (rhondasommer):

so wouldnt we just simplify it?

RhondaSommer (rhondasommer):

0.95*B=0.54*24 0.95*B=12.96 B=13.6 why would we divide both sides by sin(108)?

RhondaSommer (rhondasommer):

whoops i messed up one sec

OpenStudy (sweetburger):

well the way you did it, it would be unecessary to divide both sides by sin(108) because instead you didvided both sides by .95

OpenStudy (sweetburger):

which is basically the same thing written 2 different ways

RhondaSommer (rhondasommer):

0.951*B= 0.545*24 0.951*B=13.08 B=13.75 And oh i get it now :)

OpenStudy (sweetburger):

looks good gj @RhondaSommer

RhondaSommer (rhondasommer):

ok so now i have to find c

RhondaSommer (rhondasommer):

But ill just tag you again :)

OpenStudy (sweetburger):

ok